# Steering Language Model to Stable Speech Emotion Recognition via Contextual Perception and Chain of Thought

### v1

@@ Appendix F Effectiveness of CoT training
[…]
To validate the effectiveness of the CoT training of C2SER, we use the exact text instructions on Qwen2-Audio to conduct CoT inference. The comparison results are shown in Table 6. Obviously, Qwen2-Audio is not well capable of CoT reasoning, whose performance is significantly lower than that of C2SER. This result reveals that our CoT training boosts the reasoning capabilities of ALMs, enabling them to incorporate speech context for more accurate emotion recognition. Furthermore, Qwen2-Audio with implicit CoT performs worse than with explicit CoT, as it suffers from severe hallucinations that generate irrelevant results. This suggests that without explicit CoT training, implicit CoT fails to effectively guide reasoning and emotion recognition.

### v2

@@ VI-B Evaluation of C2SER
TABLE VIII: Ablation study of C2SER on the Emo-Emilia test set.
Model UA(%)↑ WA(%)↑ F1(%)↑
C2SER 69.00 69.00 61.61
w/o Whisper encoder 32.07 32.07 34.56
w/o Emotion2Vec-S 57.93 57.93 51.10
w/o CoT 43.14 43.14 36.15
TABLE IX: Performance improvement of C2SER after fine-tuning on MELD dataset.
Fine-tuning Epoch UA(%) ↑ WA(%) ↑ F1(%) ↑
C2SER (0 epoch) 38.66 53.10 33.10
3 43.50 58.90 38.70
6 49.30 64.86 44.10
VI-C Category Accuracy of Emotion2Vec-S and C2SER
Figure 7: Category Accuracy (%) of C2SER on the Emo-Emilia test set.
VI-D Effectiveness of CoT training
To validate the effectiveness of the CoT training of C2SER, we use the exact text instructions on Qwen2-Audio to conduct CoT inference. The comparison results are shown in Table VII. Obviously, Qwen2-Audio is not well capable of CoT reasoning, whose performance is significantly lower than that of C2SER. This result reveals that our CoT training boosts the reasoning capabilities of ALMs, enabling them to incorporate speech context for more accurate emotion recognition. Furthermore, Qwen2-Audio with implicit CoT performs worse than with explicit CoT, as it suffers from severe hallucinations that generate irrelevant results. This suggests that without explicit CoT training, implicit CoT fails to effectively guide reasoning and emotion recognition.
Having established the effectiveness of our training, we now analyze the specific advantages of our final Implicit CoT model. While its performance improvement over the Explicit CoT model is modest on the high-quality EMO-EMILIA test set (as seen in Table VII), its primary value is demonstrated in its enhanced robustness, generalization, and efficiency. The improved robustness is empirically validated by the results in Table VI, where Implicit CoT significantly outperforms Explicit CoT on more challenging and diverse datasets like M3ED and CASIA. This is because compressing the multi-step reasoning into a single, holistic step mitigates the risk of error propagation. Finally, the efficiency gain is inherent to the autoregressive process: Explicit CoT requires generating a long rationale (often >40>40 tokens), whereas Implicit CoT produces a concise expression (<10<10 tokens), guaranteeing an order-of-magnitude reduction in latency and making it viable for practical deployment.
VI-E Ablation Study
We conduct an ablation study to evaluate the contribution of each component in C2SER. The experimental results are presented in Table VIII. Firstly, removing the Whisper encoder leads to a significant degradation in performance, with the model failing to converge during explicit CoT training due to the lack of semantic perception. Secondly, the model incorporating Emotion2Vec-S outperforms the version without it, demonstrating that acoustic perception is crucial for capturing emotional expressions effectively. Finally, excluding CoT causes a substantial drop in performance. This result suggests the reasoning capability of C2SER is improved after CoT training, which leads to a more accurate and stable emotion recognition.
In addition to evaluating each component, we further examine how C2SER performs when fine-tuned on a specific dataset (MELD). As shown in Table IX, even a few epochs of fine-tuning can significantly boost the model’s performance on the target domain. However, since our primary objective is to ensure robust generalization across diverse emotional speech scenarios, we do not apply such dataset-specific fine-tuning in our main experiments.
VII Conclusion and Future Work
Despite these advances, several key challenges open avenues for future research. A primary direction is to further enhance the model’s generalization and real-world robustness. This involves not only curating more balanced and diverse training corpora to address performance variability, but also enriching the model’s contextual understanding through multimodal inputs (e.g., visual cues) and systematically evaluating its sensitivity to interactive factors like prompt phrasing. Another crucial challenge is bridging the gap between advanced reasoning capabilities and deployment efficiency. While our 7B-parameter model with implicit CoT demonstrates strong performance, future work should explore model compression and specialized fine-tuning strategies. The goal is to create more efficient variants that strike a better balance between task-specific expertise and the model’s inherent general-purpose language understanding, ultimately advancing the practical application of robust SER systems.